OpenStudy (anonymous):

In how many ways can a teacher divide a class of 21 students into groups of 7 students each?

OpenStudy (experimentx):

\[ \binom{21}{7} \times \binom{14}{7} \]

OpenStudy (anonymous):

no i dont think so the answer is suppose 66 512 160

OpenStudy (sirm3d):

\[\frac{21!}{7! \cdot 7! \cdot 7! \cdot 3!}\]

OpenStudy (experimentx):

looks like i forgot to remove are arrangements. \[ \frac{\binom{21}{7} \times \binom{14}{7} }{3!} \]